# Predicting Maintenance Cessation of Open Source Software Repositories with An Integrated Feature Framework

### v1

@@ 5. Real-World Deployment
In this section, we evaluate the practical utility of our maintenance cessation prediction system by deploying it within the openEuler ecosystem, a production-grade, community-driven Linux distribution, demonstrating both methodological generalizability and real-world impact.
@@ 5.1. Practical Risk Prediction via GBSA Classifier
To address the needs of stakeholders requiring binary, near-term risk assessments—for example, “Will this package likely cease maintenance within the next 6 months?”—we employ a Gradient Boosted Survival Analysis (GBSA) classifier using the feature set and data pipeline presented earlier. Unlike traditional regression-based survival models, GBSA excels at handling high-dimensional, time-varying, and right-censored datasets in a classification setting, and is well suited for integration into real-time package quality monitoring systems.
@@
For deployment, we treat maintenance cessation prediction as a temporal binary classification problem. Let $T$ denote a reference date (in this study, July 1, 2018), and $\Delta t$ the prediction horizon (set to 6 months). For each repository, we aggregate all features up to $T$, and label as “positive” those repositories for which unambiguous maintenance cessation occurs between $T$ and $T+\Delta t$. Repositories not ceasing maintenance in this window are used as “negative” or censored samples. We select July 1, 2018 as the reference date $T$ because a relatively large number of openEuler ecosystem packages experienced maintenance cessation during this period of time. This maximizes the number and diversity of positive samples for robust model evaluation.
@@
Our GBSA classifier achieves an overall prediction accuracy of 78.91% on an unbalanced held-out test set, confirming that our engineered features and survival-based modeling retain high discriminative power in a practical early warning context.
@@ 5.2. Deployment in the openEuler Ecosystem
openEuler (Zhou et al., 2022; openEuler, 2021) is a major open-source, Linux-based operating system, incubated within the OpenAtom Foundation and serving as a backbone for digital infrastructure in industries including Internet, finance, and telecommunications. In 2024 alone, openEuler registered over 5 million new installations (openEuler, 2024), reflecting significant exposure to the risks associated with unanticipated package maintenance cessation.
@@ 5.2.1. Mapping openEuler Packages to Upstream GitHub Repositories
openEuler is an open-source operating system whose source code is primarily hosted on the Gitee (Gitee, 2013), comprising two key repositories: the code repository for upstream source projects (openEuler, 2025a), and the package repository for build-ready software packages used in actual releases (openEuler, 2025c). These packages are directly tied to system distributions and user security. To assess ecosystem-level cessation risks, we take the openEuler release as a representative case and systematically map all its packages to their corresponding upstream GitHub repositories. Package-level metadata is retrieved from the official repository index (openEuler, 2025b).
@@
We design a seven-stage pipeline to associate openEuler-22.03-LTS aarch64/loongarch64/x86_64 packages with their corresponding upstream GitHub repositories:
  • Rule-based mapping: Manually defined rules for common packages (e.g., texlive).
  • Gitee metadata: Extracting GitHub URLs from the url field in primary.xml.
  • Spec file parsing: Locating source repositories and extracting Source0 from .spec files via depchase (fedora-modularity et al., 2017).
  • YAML field matching: Parsing git_url or src_repo fields with typo tolerance.
  • URL-based inference: Inferring likely GitHub repos from homepage domains.
  • GitHub API search: Fallback search by name, ranking by stars ($\geq$32).
  • Homepage crawling: Scraping official websites to discover embedded GitHub links.
This automated pipeline successfully maps 14,284 out of 16,888 openEuler packages (coverage: 84.6%) to valid GitHub repositories, as visualized in Table 4. The resulting mapping forms the foundation for downstream risk inference.
@@
5.2.2. Maintenance Cessation Prediction: Experimental Protocol
For each mapped upstream repository, we extract all temporal features in Section 4.1, up to the reference time $T$, and apply the trained GBSA classifier to estimate the probability of maintenance cessation within the next 6 months.
@@ 5.2.2. Maintenance Cessation Prediction: Experimental Protocol
[…]
@@ Results.
Our deployed GBSA model successfully flags 3 out of 7 repositories that truly ceased maintenance within the package set of openEuler, based on a reference date of July 1, 2018 and a 6-month prediction window. This corresponds to multiple packages, with accuracy, precision, and C-index achieving 0.9857, 1.0, and 0.8049, respectively.
@@
A manual review of the four missed cases (false negatives) shows that these repositories lacked the typical warning signals captured by our features prior to cessation; most ceased maintenance abruptly or without clear early indicators. This highlights the limitation of behavioral features in detecting entirely unannounced or exogenous cessation events. Nevertheless, our model achieves strong overall accuracy and precision, reliably identifying early risk in the vast majority of practical cases, which supports deployment in real-world package management scenarios.
@@ 6. Threats to Validity
First, our definition of maintenance cessation, relying on explicit archival status or semantic cues in documentation, may still incompletely capture all true cessation events. Silent cessations, especially among small or low-profile projects, could be missed, while some ambiguous or template-based declarations might lead to mislabeling. Additionally, although our feature engineering seeks to comprehensively model user, maintainer, and project evolution dynamics, certain abrupt or exogenous cessation events (e.g., organizational, legal, or funding shocks) may not manifest detectable signals in repository behavior or metadata prior to cessation, potentially resulting in false negatives as observed in some openEuler cases (see Section 5.2). While we perform rigorous manual validation and sensitivity analyses, latent biases in label quality and feature coverage cannot be fully excluded.
@@
Our study primarily targets GitHub-hosted repositories, which, while representing the majority of global OSS activity, may not generalize to other hosting platforms such as Gitee, GitLab, or self-hosted infrastructures. These alternatives often differ in developer conventions, community dynamics, and the richness or accessibility of behavioral data.
@@ 7. Conclusion
Maintenance cessation of open source software repositories poses major risks to software supply chains and ecosystem stability. In this work, we define a precise, actionable criterion for maintenance cessation, addressing the ambiguities of prior research. Building on this definition, we construct a comprehensive longitudinal dataset, supporting robust empirical analysis at scale. Our multi-perspective feature framework—integrating user influence, maintainer activity signals, and project evolution indicators—goes beyond traditional surface features, enabling more accurate prediction of maintenance cessation events. Using survival analysis, we achieve a high C-index (0.846), and confirmed the complementary value of our proposed features in different ways.
Finally, deployment within the openEuler ecosystem demonstrates the practical utility of our approach for proactive risk identification in real-world package management scenarios. Our interpretable modeling facilitates actionable interventions for OSS stakeholders, and paves the way for next-generation software supply chain risk management. As dependency on OSS continues to increase, scalable, early-warning tools such as ours are essential for ensuring the sustainability and security of the software ecosystem.

### v2

@@ 5. Real-World Deployment
In this section, we evaluate the practical utility of our prediction method by deploying it within the openEuler ecosystem, a production-grade, community-driven Linux distribution, demonstrating methodological generalizability and real-world impact.
@@ 5.1. Practical Risk Prediction via GBSA Classifier
To address the needs of stakeholders requiring binary, near-term risk assessments—for example, “Will this package likely cease maintenance within the next 6 months?”—we employ a Gradient Boosted Survival Analysis (GBSA) classifier using the feature set presented in Section 4. Unlike traditional regression-based survival models, GBSA excels at handling high-dimensional, time-varying, and right-censored datasets in a classification setting, and is well suited for integration into real-time package quality monitoring systems.
@@
For deployment, we treat abandonment prediction as a temporal binary classification problem. Let $T$ denote a reference date (in this study, July 1, 2018), and $\Delta t$ the prediction horizon (set to 6 months). For each repository, we aggregate all features up to $T$, and label as “positive” those repositories for which unambiguous abandonment occurs between $T$ and $T+\Delta t$. Repositories not ceasing maintenance in this window are used as “negative” or censored samples. We select July 1, 2018 as the reference date $T$ because a relatively large number of openEuler ecosystem packages experienced abandonment during this period of time. This maximizes the number and diversity of positive samples for robust model evaluation.
@@
Our GBSA classifier achieves an overall prediction accuracy of 78.91% on an unbalanced test set, confirming that our engineered features and survival-based modeling retain high discriminative power in a practical early warning context.
@@ 5.2. Deployment in the openEuler Ecosystem
openEuler (Zhou et al., 2022; openEuler, 2021) is a widely used open-source Linux OS serving critical infrastructure, with over 5 million new installations in 2024 (openEuler, 2024). This scale makes proactive package abandonment risk monitoring essential.
@@ 5.2.1. Mapping openEuler Packages to Upstream GitHub Repositories
openEuler is an open-source operating system whose source code is primarily hosted on Gitee (Gitee, 2013), comprising two key repositories: the code repository for upstream source projects (openEuler, 2025a), and the package repository for build-ready software packages used in actual releases (openEuler, 2025c). These packages are directly tied to system distributions and user security. To assess ecosystem-level abandonment risks, we take the openEuler release as a representative case and systematically map all its packages to their corresponding upstream GitHub repositories. Package-level metadata is retrieved from the official repository index (openEuler, 2025b).
@@
We design a seven-stage pipeline to associate openEuler-22.03-LTS aarch64/loongarch64/x86_64 packages with their corresponding upstream GitHub repositories: (1) Rule-based mapping: Manually defined rules for common packages (e.g., texlive); (2) Gitee metadata: Extracting GitHub URLs from the url field in primary.xml; (3) Spec file parsing: Locating source repositories and extracting Source0 from .spec files via depchase (fedora-modularity et al., 2017); (4) YAML field matching: Parsing the git_url or src_repo fields with typo tolerance; (5) URL-based inference: Inferring likely GitHub repos from homepage domains; (6) GitHub API search: Fallback search by name, ranking by stars ($\geq$32); (7) Homepage crawling: Scraping official websites to discover embedded GitHub links.
This automated pipeline successfully maps 14,284 out of 16,888 openEuler packages (coverage: 84.6%) to valid GitHub repositories, as show in Table 4. The resulting mapping forms the foundation for downstream risk inference.
@@
5.2.2. Abandonment Prediction: Experimental Protocol
For each mapped upstream repository, we extract all temporal features in Section 4.1, up to the reference time $T$, and apply the trained GBSA classifier to estimate the probability of abandonment within the next 6 months.
@@ 5.2.2. Abandonment Prediction: Experimental Protocol
We anchor prediction at $T=\text{July 1, 2018}$ and label abandonment outcomes occurring before January 1, 2019. Of the mapped repositories, 278 are included in our annotated dataset and have sufficient history and feature coverage for risk scoring. Of these, 7 repositories were actually abandoned between July 1, 2018, and January 1, 2019, forming the primary evaluation cohort.
@@ Results.
Our deployed GBSA model successfully flags 3 out of 7 repositories that were actually abandoned within the package set of openEuler, based on a reference date of July 1, 2018 and a 6-month prediction window. Accuracy, precision, and C-index achieve 0.9857, 1.0, and 0.8049, respectively. The deployment within the openEuler ecosystem further showcases how actionable and interpretable risk signals foster improved program comprehension and risk awareness throughout large-scale software supply chains.
@@
A manual review of the four missed cases (false negatives) shows that these repositories lacked the typical warning signals captured by our features prior to abandonment; most were abandoned abruptly or without clear early indicators. This highlights the limitation of behavioral features in detecting entirely unannounced or exogenous abandonment events. Nevertheless, our model achieves strong overall accuracy and precision, reliably identifying early risk in the vast majority of practical cases, which supports deployment in real-world package management scenarios.
@@ 6. Threats to Validity
First, our definition of abandonment, relying on explicit archival status or semantic cues in documentation, may still incompletely capture all true abandonment events. Silent abandonments, especially among small or low-profile projects, could be missed. Additionally, although our feature engineering seeks to comprehensively model user, maintainer, and project evolution dynamics, certain abrupt or exogenous abandonment events (e.g., organizational, legal, or funding shocks) may not manifest as detectable signals in repository behavior or metadata prior to abandonment, potentially resulting in false negatives as observed in some openEuler cases (see Section 5.2).
@@
Our study primarily targets GitHub-hosted repositories, which, while representing the majority of global OSS activity, may not generalize to other hosting platforms such as Gitee, GitLab, or self-hosted infrastructures. These alternatives may differ in developer conventions, community dynamics, and the richness or accessibility of behavioral data.
@@ 7. Conclusion
OSS project abandonment poses significant risks to software supply chains and ecosystem stability. This work introduces a precise, actionable discrimination method for abandonment, effectively addressing the low efficiency and high false-positive rates prevalent in prior studies. We construct a comprehensive longitudinal dataset to support robust empirical analysis and propose a multi-perspective feature framework that integrates user influence, maintainer activity, and project evolution indicators. Our approach significantly outperforms traditional surface features, achieving a C-index of 0.846 through survival analysis. Finally, deployment within the openEuler ecosystem demonstrates the practical utility of our approach for proactive risk identification in real-world package management scenarios.
As dependency on OSS continues to increase, scalable, early-warning tools such as ours are essential for ensuring the sustainability and security of the software ecosystem. Our interpretable framework supports program comprehension by providing actionable risk signals for dependency vetting, onboarding, and architectural planning, thus helping ensure sustainable OSS evolution.